Shows a uniform beam of mass m pivoted at its lowe
Answer #1

The arrangement shown is in static equilibrium.
If the spring is stretched by d, the force on the spring is kd
Take L as the length of the beam
a)
Equating the torque about the pivot,
L/2 x mgcos\theta = L x kd sin\theta
d = mg cot\theta / 2k

b)
Take the force exerted by the pivot be Fx and Fy
If we now equate the forces,
Fx = kd = 0.5 mg cot\theta
Fy = mg